A tank contains 18 gallons of water and is being filled at a rate of 5 gallons per minute. A second tank contains 7 gallons of water and is being filled at a rate of 6 gallons per minute. When will the two tanks contain the same amount of water?
Answer by ptaylor(2198)

Let x=number of minutes that is required for the two tanks to contain the same amount of water
In x minutes, the first tank contains 18+5x gallons
In x minutes, the second tank contains 7+6x gallons
Now, when the above two expressions are equal, the tanks will contain the same amount of water, so:
18+5x=7+6x subtract 5x and also 7 from each side
18-7+5x-5x=7-7+6x-5x collect like terms
11=x-----------------it takes 11 minutes for the two tanks to contain the same amount of water
CK
18+5*11=7+6*11
18+55=7+66
73=73
Hope this helps---ptaylor